OpenStudy (anonymous):

The half life of 14C is 5730 years. 11,460 years after an organism dies, what percent of the initial 14C will still be present in the organism's bones?

OpenStudy (frostbite):

we are going to write the radioactive decay equation as: \[N(t)=N_{0}*e ^{-k*t}\] Where \[k=\frac{ \ln(2) }{ \tau }\] Here is: τ: the half life N(t): the number of atom cores. N0: the number of atom cores to the t=0 t: the time We start by calculating k: \[k=\frac{ \ln(2) }{ 5730 }=1.210*10^{-4}\] Now as we do not know N0 from the question we can remove it by using percentage called x to the time t so we say that: \[N_{i}=x*n(y)\] Doing so we can rewrite the equation: \[x*N _{0}=N _{0}*e ^{-k*t}\] We see that N0 goes out with each other: \[x=e ^{k*t}\] We insert the information about t and k and calculation x in the following equation: \[x=e ^{-1.210*10^{-4}*11460}=0.25\] In other words there are 25% left of the 14C left.

OpenStudy (frostbite):

Correction in equation 4: \[N _{i}=x*N_{0}\]

OpenStudy (frostbite):

Sorry I forgot all the units.. but this math thingy is not that unit friendly.